# The influence of Parker spiral on the reflection-driven turbulence

### v1

@@ 1 Introduction
The solar corona and the solar wind exhibit temperatures and outflow speeds well in excess of what static, radiative-equilibrium models predict; the mechanisms that supply the required energy remain debated (Parker1958; DeMoortel2015; Cranmer2017; Chandran2021). A prominent idea is that motions in the photosphere launch Alfvénic waves into open flux tubes. These waves travel outward, and somehow dissipate, heating the plasma (Cranmer2009; DePontieu2007). This heating not only raises plasma temperature but also increases pressure, thereby contributing to wind acceleration (Tu1987; Cranmer2007). Observational support comes from measurements of ubiquitous Alfvénic fluctuations in fast solar wind streams, as reported by early spacecraft (Belcher1971; Tu1995) and consolidated in later reviews (Bruno2013). Recent in-situ data indicate that large-amplitude Alfvén waves carry sufficient energy to account for much of the required coronal heating and wind acceleration power (Halekas2023; Rivera2024).
@@
However, a fundamental theoretical limitation arises because in a homogeneous medium, Alfvén waves propagating in a single direction do not interact nonlinearly and thus cannot by themselves sustain a turbulent cascade that dissipate their energy (Kraichnan1965; Barnes1974). In the solar wind these counter-propagating fluctuations can be generated by partial reflection of outward waves from background Alfvén-speed gradients, a process termed “reflection-driven turbulence” (RDT) (Velli1989; Velli1993; Matthaeus1999; Verdini2007). Numerical and analytical models suggest this mechanism can supply a significant fraction of the heating necessary to maintain fast solar wind streams (Verdini2009; Chandran2009; Chandran2019; Chandran2021).
@@
Most models of solar wind turbulence, and RDT specifically, assume a purely radial background magnetic field for simplicity. Beyond the Alfvén point, where the solar wind speed equals the Alfvén speed, the Sun’s rotation twists the field into the well-known Parker spiral (PS) (Parker1958; Weber1967). Close to the Sun, the spiral angle is very small, but beyond approximately 0.30.3-11AU the azimuthal component becomes significant, such that it could alter wave propagation and nonlinear coupling (Verdini2008; Owens2013; Verdini2018a).
@@
The goal of this paper is to quantify how PS geometry modifies reflection-driven turbulence and the heating it causes in the super-Alfvénic solar wind. We argue that the relevant change to the standard Dmitruk2002 phenomenology of RDT due to the PS is that the evolution of the characteristic scales of the turbulence in the perpendicular and parallel directions are modified by the mean field’s direction. We then study this phenomenology using the expanding-box model (EBM) (Grappin1993) to follow a small plasma parcel advected outward, presenting a suite of three-dimensional compressible MHD simulations initialized with strongly outward-dominated $\bm{z}^{+}$ fluctuations. We find that the RDT picture remains broadly valid in PS geometry—the same reflection-driven dynamics operate, but the controlling scales change. Perpendicular expansion stretches eddies into pancake-like structures, while the growing azimuthal field in the PS rotates the mean field relative to those pancakes and therefore “cuts across” them; this geometric change produces smaller effective perpendicular scales than in the radial case. As a result the two transverse correlation lengths evolve differently and the effective perpendicular outer scale in a PS background tends to saturate rather than grow indefinitely. This causes the expansion-to-nonlinearity ratio, termed $\chi_{\rm exp}$, to decrease less rapidly and the system remains in a sustained, imbalanced nonlinear state. By contrast, in radial geometry the perpendicular scale grows with expansion, $\chi_{\rm exp}$ falls, and the cascade can shut off, leaving a magnetically dominated state with little associated heating.
@@
The paper is organized as follows. In section˜2 we develop the theoretical framework for turbulence evolution and introduce the compressible expanding-box magnetohydrodynamic (MHD) model employed in our study. We extend the simple reflection-driven phenomenology of Dmitruk2002 and show how PS geometry alters reflection, projected perpendicular scales, and the resulting nonlinear time-scales. In section˜3 we summarize the numerical method, list key parameters and resolution choices, and specify the initial conditions used in the simulation suite. In section˜4, we test the theoretical expectations of section˜2 by examining evolution of outward/inward energies, normalized cross-helicity, and the competing nonlinear and expansion timescales demonstrating how these diagnostics depend on the PS in general agreement with the phenomenology. Section˜5 explores various other diagnostics for the purpose of comparing to observations, reporting the parametric evolution of cross-helicity and residual energy, switchback statistics, magnetic compressibility, and synthetic fly-by traces for comparison with spacecraft measurements. Finally, section˜6 summarizes the main results, discusses their implications for existing theory and in-situ observations.
@@ 2 Theoretical Framework
In this section, we build a theoretical model to describe how solar-wind expansion and large-scale magnetic field geometry shape Alfvénic turbulence. We introduce the expanding-box formalism to capture the kinematic effects of transverse expansion on fields and their length scales (Grappin1993), and formulate the dynamics in terms of Elsässer variables to make the reflection-driven coupling explicit. Building on this foundation, we outline reflection-driven turbulence and derive the Dmitruk2002 decay phenomenology, extending this to include the Parker spiral. This framework both motivates the simulations and diagnostics developed later and provides testable predictions for our simulations.
@@ 2.1 Governing Equations for Reflection-Driven Turbulence
The solar wind’s turbulent evolution is governed by the interplay between reflection and nonlinear interactions. To model this in regions beyond the Alfvén critical point, where the flow becomes super-Alfvénic, we adopt the Expanding Box Model (EBM; Grappin1993), which approximates the spherical expansion of the solar wind in a Cartesian patch of wind co-moving with the constant radial flow velocity $U$. Aligning the $x$-axis with the radial direction, the EBM introduces an expansion factor $a(t)=R(t)/R_{0}$, where $R(t)=R_{0}+Ut$ is heliocentric distance for some initial $R_{0}$, and modifies spatial gradients as $\tilde{\nabla}=(\partial_{x},a^{-1}\partial_{y},a^{-1}\partial_{z})$, viz., x{x} (radial) is unstretched, and y,z{y},{z} (azimuthal/normal) are stretched. $\dot{a}=U/R_{0}$ is the constant expansion rate due to $U$. The governing MHD equations in this frame become (Grappin1993):
@@
$\frac{\partial\bm{u}}{\partial t}+(\bm{u}\cdot\tilde{\nabla})\bm{u}$ $=-\frac{\tilde{\nabla}p}{\rho}+\frac{(\bm{B}\cdot\tilde{\nabla})\bm{B}}{4\pi\rho}-\frac{\dot{a}}{a}\mathbb{T}\cdot\bm{u},$ (2)
$\frac{\partial\bm{B}}{\partial t}+(\bm{u}\cdot\tilde{\nabla})\bm{B}$ $=(\bm{B}\cdot\tilde{\nabla})\bm{u}-\frac{\dot{a}}{a}\mathbb{L}\cdot\bm{B},$ (3)
where $\mathbb{T}=\text{diag}(0,1,1)$ and $\mathbb{L}=\text{diag}(2,1,1)$ encode anisotropic damping from angular momentum conservation and magnetic flux freezing, respectively. We assume a locally isothermal equation of state, $p=c_{s}^{2}\rho$, where $p$ is the thermal pressure, $c_{s}$ is the sound speed, and $\rho$ is the plasma density. $c_{s}$ changes with $a$ as $c_{s}\propto a^{-2/3}$, representing the adiabatic cooling of the solar wind with expansion.
@@
$\frac{\partial\bm{z}^{\pm}}{\partial t}\pm\left(\bm{v}_{A}\cdot\tilde{\nabla}\right)\bm{z}^{\pm}+\left(\bm{z}^{\mp}\cdot\tilde{\nabla}\right)\bm{z}^{\pm}+\tilde{\nabla}p^{\prime\pm}=-\frac{\dot{a}}{2a}\mathbb{T}\cdot\left(\bm{z}^{+}+\bm{z}^{-}\right)-\frac{\dot{a}}{2a}\left(z_{x}^{\pm}-z_{x}^{\mp}\right)\hat{\bm{x}}$, (7)
where $\tilde{\nabla}p^{\prime\pm}$ ensures $\nabla\cdot\bm{z}^{\pm}=0$. We take the mean magnetic field $\bar{\bm{B}}$ to point in the positive radial direction, so that $\bm{z}^{+}$ perturbations propagate outward in the absence of reflection.
@@ 2.2 Wave-Action Elsässer Variables
In an expanding medium like the solar wind, Alfvén wave amplitudes can naturally decay due to geometric WKB effects as well as true dissipation (Heinemann1980). To remove this trivial evolution, which manifests in Eq. (7) via $\bm{z}^{\pm}$ terms on right hand side, we rescale the Elsässer variables as (Chandran2009; Meyrand2025),
@@
$\dot{a}\frac{\partial\tilde{\bm{z}}^{\pm}}{\partial a}\pm\left(\bm{v}_{A}\cdot\tilde{\nabla}\right)\tilde{\bm{z}}^{\pm}+{a^{-1/2}}\left(\tilde{\bm{z}}^{\mp}\cdot\tilde{\nabla}\right)\tilde{\bm{z}}^{\pm}+\tilde{\nabla}\tilde{p}^{\prime}=-\frac{\dot{a}}{2a}\tilde{\mathbb{T}}\cdot\tilde{\bm{z}}^{\mp}$, (9)
@@ 2.3 Energy Dynamics in Expanding Solar Wind Turbulence
$\dot{a}\frac{\partial}{\partial a}{\tilde{E}^{\pm}}=-\frac{\dot{a}}{4a}{\langle\tilde{\bm{z}}^{+}\cdot\mathbb{T}\cdot\tilde{\bm{z}}^{-}\rangle}+\frac{\dot{a}}{4a}\langle\tilde{z}_{x}^{+}\tilde{z}_{x}^{-}\rangle-\bm{D}^{\pm},$ (10)
here $\bm{D}^{\pm}$ is dissipation term due to the viscosity/resistivity in the Eqs. (2.1)-(2.3). Because we expect predominantly Alfvénic fluctuations, which are perpendicular to $\bar{\bm{B}}$, in the case of the purely radial field the term $\langle\tilde{z}_{x}^{+}\tilde{z}_{x}^{-}\rangle$ is likely negligible, so the energy evolution simplifies to $\dot{a}\,\partial\tilde{E}/\partial a=-(\dot{a}/2a)\tilde{E}_{r}$ (Meyrand2025). The reflection thus acts as a source or sink of wave-action energy depending on the sign of the residual energy $\tilde{E}_{r}$ (note that any change in fluctuation energy is balanced by the background-flow energy budget; Chandran_2015). In the Parker spiral configuration, the $\langle\tilde{z}_{x}^{+}\tilde{z}_{x}^{-}\rangle$ term could become significant even when $\bm{z}^{\pm}$ is predominantly perpendicular to $\bm{B}$, and we likewise see from Eq. (9) that it is naturally driven with the correlation required to act as a source of $\tilde{E}^{\pm}$.
@@ 2.4 Reflection-driven Turbulence Phenomenology
where $\mathcal{L}^{\pm}=\bm{v}_{A}\cdot\tilde{\nabla}$ represents Alfvénic propagation, $\mathcal{N}^{\pm}={\bm{z}}^{\mp}\cdot\tilde{\nabla}=a^{-1/2}\tilde{\bm{z}}^{\mp}\cdot\tilde{\nabla}\sim{z}^{\mp}/\ell_{\perp}$ represents nonlinear terms that will damp out ${z}^{\pm}$ (here $\ell_{\perp}$ denotes the perpendicular correlation length scale of fluctuations and ${z}^{+}$ denotes the characteristic outer-scale amplitude of outward-propagating fluctuations at scale $\ell_{\perp}$), and reflection is $\mathcal{R}^{\pm}=-\dot{a}/2a\mathbb{\tilde{T}}$. The characteristic timescales associated with the terms in Eq. (11) can differ in their scaling. The Alfvén propagation time along the mean field is $\tau_{\rm A}=\ell_{\parallel}/v_{\rm A}$ ($\ell_{\parallel}$ is the parallel correlation length), while the nonlinear interaction time across the field is $\tau_{\rm nl}=\ell_{\perp}/{z}^{\mp}$. Both $\tau_{\rm A}$ and $\tau_{\rm nl}$ decrease at smaller scales, since they depend directly on the characteristic sizes of turbulent structures along and across the mean magnetic field respectively. In contrast, the expansion time $\tau_{\rm exp}=a/\dot{a}$ is independent of scale. Following Meyrand2025, we define the dimensionless ratios;
@@
which quantify the relative importance of expansion and Alfvén propagation compared to nonlinear interactions (GS95). These ratios will play a key role in controlling the dynamics of the turbulent cascade in an expanding solar wind.
The standard phenomenology (Dmitruk2002; Verdini2007; Chandran2009) balances these terms to model turbulence and heating, making two key balancing assumptions. First, for $\tilde{z}^{-}$ (backward waves) propagation and time variation are assumed negligible, the reflection source ($\mathcal{R}^{-}\tilde{z}^{+}$) balances the nonlinear sink ($\mathcal{N}^{-}\tilde{z}^{-}$):
@@
This corresponds to $z^{+}\propto a^{-1}$ or (in radial field) $z^{+}/v_{\rm A}\sim\text{const}$. The heating rate $\mathcal{Q}$, which is the rate at which energy in the outward-propagating mode $\tilde{z}^{+}$ is damped, can be calculated using Eq. (10) by multiplying with $\rho V$ to get actual energy (see, Perez2021) [^1] From Eq. 2.12 of Perez2021, neglecting the residual energy and $E^{-}$, then assuming super-Alfvénic flow with $U=\mathrm{const}$ and $\hat{\bm{b}}=\hat{\bm{r}}$ (where $\hat{\bm{r}}$ is the radial unit vector in spherical coordinates), the steady energy balance reduces to $\nabla\cdot(\hat{\bm{r}}\,U\,\rho\,E^{+})+\tfrac{1}{2}\,\rho E^{+}U\,\nabla\cdot\hat{\bm{r}}=-Q$. Using $\nabla\cdot\hat{\bm{r}}=2/r$, this becomes $\partial_{r}(r^{2}\rho UE^{+})/r^{2}+\rho UE^{+}/r=-{Q}$. With $U=\mathrm{const}$ and $\rho r^{2}=\rho_{0}$, we get $U\rho_{0}\big(\tfrac{1}{r^{2}}\partial_{r}E^{+}+\tfrac{1}{r^{3}}E^{+}\big)=-\mathcal{Q}$. Since $E^{+}=r\tilde{E}^{+}$, we obtain $U\rho_{0}\big[\tfrac{1}{r^{2}}\partial_{r}(\tilde{E}^{+}/r)+\tfrac{1}{r^{4}}\tilde{E}^{+}\big]=U\rho_{0}/{r^{3}}\partial_{r}\tilde{E}^{+}=-{Q}$. This matches Eq. (16), if we note that the extra $1/r^{2}$ arises because this is heating per unit volume, whereas Eq. (15) is heating per unit co-moving volume. , where $V\propto a^{2}$ is the volume of the domain. This gives
@@
where we neglected the residual energy and use $\langle|\tilde{\bm{z}}^{+}|\rangle^{2}\approx(\tilde{z}^{+})^{2}$ (mean square amplitude). The phenomenology in section˜2.3 predicts $\tilde{E}\propto a^{-1}$, giving $\mathcal{Q}\propto a^{-3}$. Here $\mathcal{Q}$ is the heating rate per co-moving volume (not per physical volume); this $\mathcal{Q}\propto a^{-5}$ when expressed per unit physical volume. Interestingly, this heating rate does not depend on the nonlinear rate or scale of the turbulence, because a smaller transverse scale $z^{+}$ damps $z^{-}$ more rapidly, a smaller $\ell_{\perp}$ implies that $z^{-}$ has smaller amplitude for the same reflection rate $\mathcal{R}$. This lower amplitude then cancels the increased nonlinear rate coming from the smaller scale of $z^{-}$, thus leading to the same overall efficiency at damping $z^{+}$.

### v2

@@ 1 Introduction
The solar corona and the solar wind exhibit temperature and velocity profiles that cannot be fully explained by Parker (1958)’s simple isothermal hydrodynamic model. The mechanisms that supply the energy required to explain the differences remain debated (De Moortel and Browning, 2015; Cranmer et al., 2017; Chandran, 2021). A prominent idea is that motions in the photosphere launch Alfvénic waves into open flux tubes. These waves travel outward and somehow dissipate, heating the plasma (Cranmer, 2009; De Pontieu et al., 2007), both raising plasma temperature and increasing magnetic or wave pressure, thereby contributing to wind acceleration (Tu, 1987; Cranmer et al., 2007). Observational support comes from measurements of ubiquitous Alfvénic fluctuations in fast solar wind streams, as reported by early spacecraft (Belcher and Davis, 1971; Tu and Marsch, 1995) and consolidated in later reviews (Bruno and Carbone, 2013). Recent in-situ data indicate that large-amplitude Alfvén waves can carry substantial energy for coronal heating and wind acceleration (Halekas et al., 2023; Rivera et al., 2024).
@@
However, a fundamental theoretical limitation arises because in a homogeneous medium, Alfvén waves propagating in a single direction do not interact nonlinearly and thus cannot by themselves sustain a turbulent cascade that dissipate their energy (Kraichnan, 1965; Barnes and Hollweg, 1974). In the solar wind, counter-propagating fluctuations can be generated by partial reflection of outward waves from background Alfvén-speed gradients, a process termed “reflection-driven turbulence” (RDT) (Velli et al., 1989; Velli, 1993; Matthaeus et al., 1999; Verdini and Velli, 2007). Numerical and analytical models suggest this mechanism can supply a significant fraction of the heating necessary to maintain fast solar wind streams (Verdini et al., 2009; Chandran and Hollweg, 2009; Chandran and Perez, 2019; Chandran, 2021). We note that when compressibility is included, parametric decay instability can also generate counter-propagating and compressible fluctuations from finite-amplitude Alfvén waves, though the relative importance of this mechanism in solar wind turbulence evolution remains debated (Malara et al., 2000).
@@
Most models of solar wind turbulence, and RDT specifically, assume a purely radial background magnetic field for simplicity. Beyond the Alfvén point, where the solar wind speed equals the Alfvén speed, the Sun’s rotation twists the field into the well-known Parker spiral (PS) (Parker 1958; Weber and Davis 1967). Close to the Sun, the spiral angle is very small, but beyond approximately 0.30.3-11AU the azimuthal component becomes significant, such that it could alter wave propagation and nonlinear coupling (Verdini et al., 2008; Owens and Forsyth, 2013; Verdini et al., 2018).
@@
The goal of this paper is to quantify how PS geometry modifies reflection-driven turbulence and the heating it causes in the super-Alfvénic solar wind. We argue that the relevant change to the standard Dmitruk et al. (2002) phenomenology of RDT due to the PS is that the evolution of the characteristic scales of the turbulence in the perpendicular and parallel directions are modified by the mean field’s direction. We then study this phenomenology using the expanding-box model (EBM) (Grappin et al., 1993) to follow a small plasma parcel advected outward, presenting a suite of three-dimensional compressible MHD simulations initialized with strongly outward-dominated $\bm{z}^{+}$ fluctuations. We find that the RDT picture remains broadly valid in PS geometry—the same reflection-driven dynamics operate, but the controlling scales change. Perpendicular expansion stretches eddies into pancake-like structures, while the growing azimuthal field in the PS rotates the mean field relative to those pancakes and therefore “cuts across” them; this geometric change produces smaller effective perpendicular scales than in the radial case. As a result the two transverse correlation lengths evolve differently and the effective perpendicular outer scale in a PS background tends to saturate rather than grow indefinitely. This causes the expansion-to-nonlinearity ratio, termed $\chi_{\rm exp}$, to decrease less rapidly and the system remains in a sustained, imbalanced nonlinear state. By contrast, in radial geometry the perpendicular scale grows with expansion, $\chi_{\rm exp}$ falls, and the cascade can shut off, leaving a magnetically dominated state with little associated heating.
@@
The paper is organized as follows. In section˜2 we develop the theoretical framework for turbulence evolution and introduce the compressible expanding-box magnetohydrodynamic (MHD) model employed in our study. We extend the simple reflection-driven phenomenology of Dmitruk et al. (2002) and show how PS geometry alters reflection, projected perpendicular scales, and the resulting nonlinear time-scales. In section˜3 we summarize the numerical method, list key parameters and resolution choices, and specify the initial conditions used in the simulation suite. In section˜4, we test the theoretical expectations of section˜2 by examining evolution of outward/inward energies, normalized cross-helicity, and the competing nonlinear and expansion timescales demonstrating how these diagnostics depend on the PS in general agreement with the phenomenology. Section˜5 explores various other diagnostics for the purpose of comparing to observations, reporting the parametric evolution of cross-helicity and residual energy, switchback statistics, magnetic compressibility, and synthetic fly-by traces for comparison with spacecraft measurements. Finally, section˜6 summarizes the main results, discusses their implications for existing theory and in-situ observations.
@@ 2 Theoretical Framework
In this section, we build a theoretical model to describe how solar-wind expansion and large-scale magnetic field geometry shape Alfvénic turbulence. We introduce the expanding-box formalism to capture the kinematic effects of transverse expansion on fields and their length scales (Grappin et al., 1993), and formulate the dynamics in terms of Elsässer variables to make the reflection-driven coupling explicit. Building on this foundation, we outline reflection-driven turbulence and derive the Dmitruk et al. (2002) decay phenomenology, extending this to include the Parker spiral. This framework both motivates the simulations and diagnostics developed later and provides testable predictions for our simulations.
@@ 2.1 Governing Equations for Reflection-Driven Turbulence
The solar wind’s turbulent evolution is governed by the interplay between reflection and nonlinear interactions. To model this in regions beyond the Alfvén critical point, where the flow becomes super-Alfvénic, we adopt the Expanding Box Model (EBM; Grappin et al. 1993), which approximates the spherical expansion of the solar wind in a Cartesian patch of wind co-moving with the constant radial flow velocity $U$. Aligning the $x$-axis with the radial direction, the EBM introduces an expansion factor $a(t)=R(t)/R_{0}$, where $R(t)=R_{0}+Ut$ is heliocentric distance for some initial $R_{0}$, and modifies spatial gradients as $\tilde{\nabla}=(\partial_{x},a^{-1}\partial_{y},a^{-1}\partial_{z})$, viz., x{x} (radial) is unstretched, and y,z{y},{z} (azimuthal/normal) are stretched. $\dot{a}=U/R_{0}$ is the constant expansion rate due to $U$. The governing MHD equations in this frame become:
@@
$\frac{\partial\bm{u}}{\partial t}+(\bm{u}\cdot\tilde{\nabla})\bm{u}$ $=-\frac{\tilde{\nabla}p}{\rho}+\frac{(\bm{B}\cdot\tilde{\nabla})\bm{B}}{4\pi\rho}-\frac{\dot{a}}{a}\mathbb{T}\cdot\bm{u}+\mathcal{D}_{u},$ (2)
$\frac{\partial\bm{B}}{\partial t}+(\bm{u}\cdot\tilde{\nabla})\bm{B}$ $=(\bm{B}\cdot\tilde{\nabla})\bm{u}-\frac{\dot{a}}{a}\mathbb{L}\cdot\bm{B}+\mathcal{D}_{B},$ (3)
where $\mathbb{T}=\text{diag}(0,1,1)$ and $\mathbb{L}=\text{diag}(2,1,1)$ encode anisotropic damping from angular momentum conservation and magnetic flux freezing, respectively (here diag denotes a diagonal matrix with the specified values), while $\mathcal{D}_{u}$ and $\mathcal{D}_{B}$ are the dissipative (viscous and resistive) terms that act only at small scales. In the real solar wind they correspond to kinetic damping processes, while in simulations they mainly represent explicit and numerical dissipation near the grid scale.
We assume a locally isothermal equation of state, $p=c_{s}^{2}\rho$, where $p$ is the thermal pressure, $c_{s}$ is the sound speed, and $\rho$ is the plasma density. By "locally isothermal," we mean that the temperature is spatially uniform throughout the computational domain at any given instant during the simulation. However, this does not imply that the temperature remains constant with heliocentric distance. Rather, we allow the temperature (and thus $c_{s}$) to evolve with $a$ in a manner that reproduces the cooling expected for expanding gas. In particular, the sound speed varies with the expansion as $c_{s}\propto a^{-2/3}$, representing the cooling of the solar wind as it expands. The locally isothermal approximation thus captures the background thermal evolution while simplifying the thermodynamics by enforcing spatial temperature uniformity at each time step.
@@
$\frac{\partial\bm{z}^{\pm}}{\partial t}\pm\left(\bm{v}_{A}\cdot\tilde{\nabla}\right)\bm{z}^{\pm}+\left(\bm{z}^{\mp}\cdot\tilde{\nabla}\right)\bm{z}^{\pm}+\tilde{\nabla}p^{\prime\pm}=-\frac{\dot{a}}{2a}\mathbb{T}\cdot\left(\bm{z}^{+}+\bm{z}^{-}\right)$
$-\frac{\dot{a}}{2a}\left(z_{x}^{\pm}-z_{x}^{\mp}\right)\hat{\bm{x}}+\mathcal{D}^{\pm},$ (7)
where $\mathcal{D}^{\pm}$ represents the dissipative effects due to viscosity/resistivity that act on $\bm{z}^{\pm}$ , and $\tilde{\nabla}p^{\prime\pm}$ ensures $\nabla\cdot\bm{z}^{\pm}=0$. We take the mean magnetic field $\bar{\bm{B}}$ to point in the positive radial direction, so that $\bm{z}^{+}$ perturbations propagate outward in the absence of reflection.
@@ 2.2 Wave-Action Elsässer Variables
In an expanding medium like the solar wind, Alfvén wave amplitudes can naturally decay due to geometric WKB effects as well as true dissipation (Heinemann and Olbert, 1980). To remove this trivial evolution, which manifests in Eq. (7) via $\bm{z}^{\pm}$ terms on right hand side, we rescale the Elsässer variables as (Chandran and Hollweg, 2009; Meyrand et al., 2025),
@@
$\dot{a}\frac{\partial\tilde{\bm{z}}^{\pm}}{\partial a}\pm\left(\bm{v}_{A}\cdot\tilde{\nabla}\right)\tilde{\bm{z}}^{\pm}+{a^{-1/2}}\left(\tilde{\bm{z}}^{\mp}\cdot\tilde{\nabla}\right)\tilde{\bm{z}}^{\pm}+\tilde{\nabla}\tilde{p}^{\prime}=-\frac{\dot{a}}{2a}\tilde{\mathbb{T}}\cdot\tilde{\bm{z}}^{\mp}+\mathcal{\tilde{D}}^{\pm}$, (9)
@@ 2.3 Energy Dynamics in Expanding Solar Wind Turbulence
$\dot{a}\frac{\partial}{\partial a}{\tilde{E}^{\pm}}=-\frac{\dot{a}}{4a}{\langle\tilde{\bm{z}}^{+}\cdot\mathbb{T}\cdot\tilde{\bm{z}}^{-}\rangle}+\frac{\dot{a}}{4a}\langle\tilde{z}_{x}^{+}\tilde{z}_{x}^{-}\rangle+\mathcal{\tilde{D}}^{\pm}$, (10)
because we expect predominantly Alfvénic fluctuations, which are perpendicular to $\bar{\bm{B}}$, in the case of the purely radial field the term $\langle\tilde{z}_{x}^{+}\tilde{z}_{x}^{-}\rangle$ is likely negligible, so the energy evolution simplifies to $\dot{a}\,\partial\tilde{E}/\partial a=-(\dot{a}/2a)\tilde{E}_{r}$ (Meyrand et al., 2025). The reflection thus acts as a source or sink of wave-action energy depending on the sign of the residual energy $\tilde{E}_{r}$ (note that any change in fluctuation energy is balanced by the background-flow energy budget; Chandran et al. 2015). In the absence of dissipation and external forcing, $\tilde{E}^{+}-\tilde{E}^{-}$ is conserved by the combined dynamics of wave reflection and advection, because the corresponding source terms are identical in the ++ and - equations and cancel upon subtraction; making this difference the true wave-action invariant (Heinemann and Olbert, 1980).
In the Parker spiral configuration, the $\langle\tilde{z}_{x}^{+}\tilde{z}_{x}^{-}\rangle$ term could become significant even when $\bm{z}^{\pm}$ is predominantly perpendicular to $\bm{B}$, and we likewise see from Eq. (9) that it is naturally driven with the correlation required to act as a source of $\tilde{E}^{\pm}$.
@@ 2.4 Reflection-driven Turbulence Phenomenology
where $\mathcal{L}^{\pm}=\bm{v}_{A}\cdot\tilde{\nabla}$ represents Alfvénic propagation, $\mathcal{N}^{\pm}={\bm{z}}^{\mp}\cdot\tilde{\nabla}=a^{-1/2}\tilde{\bm{z}}^{\mp}\cdot\tilde{\nabla}\sim{z}^{\mp}/\ell_{\perp}$ represents nonlinear terms that will damp out ${z}^{\pm}$ (here $\ell_{\perp}$ denotes the perpendicular correlation length scale of fluctuations and ${z}^{+}$ denotes the characteristic outer-scale amplitude of outward-propagating fluctuations at scale $\ell_{\perp}$), and reflection is $\mathcal{R}^{\pm}=-\dot{a}/2a\mathbb{\tilde{T}}$. The characteristic timescales associated with the terms in Eq. (11) can differ in their scaling. The Alfvén propagation time along the mean field is $\tau_{\rm A}=\ell_{\parallel}/v_{\rm A}$ ($\ell_{\parallel}$ is the parallel correlation length), while the nonlinear interaction time across the field is $\tau_{\rm nl}=\ell_{\perp}/{z}^{\mp}$. Both $\tau_{\rm A}$ and $\tau_{\rm nl}$ decrease at smaller scales, since they depend directly on the characteristic sizes of turbulent structures along and across the mean magnetic field respectively. In contrast, the expansion time $\tau_{\rm exp}=a/\dot{a}$ is independent of scale. Following Meyrand et al. (2025), we define the dimensionless ratios;
@@
which quantify the relative importance of expansion and Alfvén propagation compared to nonlinear interactions (Goldreich and Sridhar, 1995). These ratios will play a key role in controlling the dynamics of the turbulent cascade in an expanding solar wind.
The standard phenomenology (Dmitruk et al., 2002; Verdini and Velli, 2007; Chandran and Hollweg, 2009) balances these terms to model turbulence and heating, making two key balancing assumptions. First, for $\tilde{z}^{-}$ (backward waves) propagation and time variation are assumed negligible, the reflection source ($\mathcal{R}^{-}\tilde{z}^{+}$) balances the nonlinear sink ($\mathcal{N}^{-}\tilde{z}^{-}$):
@@
This corresponds to $z^{+}\propto a^{-1}$ or (in radial field) $z^{+}/v_{\rm A}\sim\text{const}$. The heating rate $\mathcal{Q}$, which is the rate at which energy in the outward-propagating mode $\tilde{z}^{+}$ is damped, can be calculated using Eq. (10) by multiplying with $\rho V$ to get actual energy (see, Perez et al. 2021) [^1] From Eq. 2.12 of Perez et al. (2021), neglecting the residual energy and $E^{-}$, then assuming super-Alfvénic flow with $U=\mathrm{const}$ and $\hat{\bm{b}}=\hat{\bm{r}}$ (where $\hat{\bm{r}}$ is the radial unit vector in spherical coordinates), the steady energy balance reduces to $\nabla\cdot(\hat{\bm{r}}\,U\,\rho\,E^{+})+\tfrac{1}{2}\,\rho E^{+}U\,\nabla\cdot\hat{\bm{r}}=-Q$. Using $\nabla\cdot\hat{\bm{r}}=2/r$, this becomes $\partial_{r}(r^{2}\rho UE^{+})/r^{2}+\rho UE^{+}/r=-{Q}$. With $U=\mathrm{const}$ and $\rho r^{2}=\rho_{0}$, we get $U\rho_{0}\big(\tfrac{1}{r^{2}}\partial_{r}E^{+}+\tfrac{1}{r^{3}}E^{+}\big)=-\mathcal{Q}$. Since $E^{+}=r\tilde{E}^{+}$, we obtain $U\rho_{0}\big[\tfrac{1}{r^{2}}\partial_{r}(\tilde{E}^{+}/r)+\tfrac{1}{r^{4}}\tilde{E}^{+}\big]=U\rho_{0}/{r^{3}}\partial_{r}\tilde{E}^{+}=-{Q}$. This matches Eq. (16), if we note that the extra $1/r^{2}$ arises because this is heating per unit volume, whereas Eq. (15) is heating per unit co-moving volume. , where $V\propto a^{2}$ is the volume of the domain. This gives
@@
where we neglected the residual energy and use $\langle|\tilde{\bm{z}}^{+}|\rangle^{2}\approx(\tilde{z}^{+})^{2}$ (mean square amplitude). The phenomenology in section˜2.3 predicts $\tilde{E}\propto a^{-1}$, which implies a heating rate per co-moving volume $\mathcal{Q}\propto a^{-3}$. Converting to physical units (physical volume $\propto a^{2}$) gives a heating rate per physical volume $\mathcal{\tilde{Q}}\propto a^{-5}$. The heating rate $\mathcal{Q}$ represents the actual dissipation of turbulent energy into thermal energy through viscous and resistive processes, not the reversible redistribution of energy between outward and inward propagating modes. Reflection converts $\bm{z}^{+}$ into $\bm{z}^{-}$ and vice versa, and the nonlinear turbulent cascade transfers energy from large scales to small scales, but only when the cascade reaches dissipative scales is energy irreversibly removed and counted in $\mathcal{Q}$. In our simulations, explicit viscosity and resistivity are not included; instead, numerical dissipation at the grid scale in Athena++ serves as an effective dissipation mechanism that captures the cascade. The rate $\mathcal{Q}$ therefore measures how rapidly the outward energy $\tilde{E}^{+}$ is depleted by this dissipative process, which is fundamentally distinct from the reversible energy exchange between $\bm{z}^{+}$ and $\bm{z}^{-}$ driven by reflection. We note that while kinetic effects likely dominate dissipation in the actual solar wind plasma, our MHD framework with numerical dissipation at the grid scale captures the essential phenomenology of energy removal at small scales.
@@
Interestingly, this heating rate does not depend on the nonlinear rate or scale of the turbulence, because a smaller transverse scale $z^{+}$ damps $z^{-}$ more rapidly, a smaller $\ell_{\perp}$ implies that $z^{-}$ has smaller amplitude for the same reflection rate $\mathcal{R}$. This lower amplitude then cancels the increased nonlinear rate coming from the smaller scale of $z^{-}$, thus leading to the same overall efficiency at damping $z^{+}$.